1.   What is the momentum of a proton traveling at 𝑣 = 0.68𝑐?



1
Expert's answer
2022-04-04T14:43:34-0400

Rest energy of proton is "mc^2=938 MeV"

Speed of proton is "v=0.68c"

The expression for the relativistic momentum of proton is

"p=\frac{1}{\sqrt(1-(\frac{v}{c})^2)} mv"

"=\frac{1}{\sqrt(1-(\frac{0.68c}{c})^2)} (938 MeV/c^2)(0.68 c)"

"=869.9 MeV\/c"

=870 MeV/c